# : A Quantum Multiple Kernel Learning Approach for Drug Discovery

### v1

@@ II-C Machine Learning Enhancement and Quantum Frontiers
Quantitative Structure-Activity Relationship () models aim to establish a quantitative link between the chemical structure of a molecule and its biological activity [5]. These mathematical models provide an in silico methodology to predict the properties of novel compounds, thereby prioritizing experimental efforts and reducing costs. Applications range from predicting pharmacokinetic processesβabsorption, distribution, metabolism, and excretion ()βto assessing toxicological endpoints, and even modelling new properties
@@ III Methodology
III-A Dataset and Preprocessing
The study utilizes a public dataset of (Dual-specificity tyrosine-phosphorylation-regulated kinase 1A) inhibitors [11].
The dataset was originally shared by a second research group, which made the model classically, so it is our starting point for performing a benchmark against the same dataset.
A comprehensive dataset of 1291 molecules exhibiting inhibitory activity against human kinase was systematically compiled from the database (). represents a critical therapeutic target for neurodegenerative disorders, particularly Alzheimerβs disease, due to its role in tau phosphorylation and neuronal dysfunction [12].
@@ III-A1 Molecular Descriptors
NPR1 (Normalized Principal Ratio 1): The ratio PMI1/PMI3, normalized to eliminate molecular size effects. NPR1 values range from 0.5 (spherical) to 1.0 (linear), providing size-independent shape characterization essential for scaffold-hopping applications
@@
MolLogP (Molecular LogP): Octanol-water partition coefficient calculated using the Wildman-Crippen method implemented in . Values typically range from -3 to +8 for drug-like compounds, with optimal range 1-3 for oral bioavailability
TPSA (Topological Polar Surface Area): Sum of surfaces of polar atoms (oxygen and nitrogen) in a molecule, calculated using group contributions. TPSA values Β‘140 Γ Β² correlate with good blood-brain barrier permeability
NumRotatableBonds: Count of rotatable bonds in the molecule, excluding amide bonds. Drug-like molecules typically have 10 rotatable bonds according to Lipinskiβs rule
@@
BertzCT (Bertz Complexity Index): A graph-theoretical descriptor that quantifies molecular complexity by considering both the number of atoms and their connectivity patterns. Typical values range from 0 for simple molecules (e.g., methane) to over 1000 for complex natural products [14]. Its considers both the number of atoms and their connectivity patterns within a molecular structure. This index helps in understanding and predicting the complexity of molecules, particularly in the context of synthetic planning and analysis of synthetic strategies. The Bertz Complexity Index utilizes graph theory to represent molecules. In this representation, atoms are nodes (vertices) and bonds are edges within the graph
Chi0n, Chi1n (Kier-Hall Connectivity Indices): Topological indices describing molecular branching and connectivity. Chi0n reflects molecular size, while Chi1n captures first-order connectivity patterns
@@
NumHBD (Hydrogen Bond Donors): Count of hydrogen bond donor groups (OH, NH). Drug-like molecules typically have 5 donors
NumHBA (Hydrogen Bond Acceptors): Count of hydrogen bond acceptor atoms (N, O). Optimal range 10 for oral drugs
FractionCsp3: Fraction of carbons that are sp3 hybridized. Higher Fsp3 values () correlate with increased three-dimensionality and improved developability
@@ Acknowledgment
This work was supported by the project funded by Ministerio de Ciencia, InnovaciΓ³n y Universidades, through .

### v3

@@ II-A Foundations of Quantitative Structure-Activity Relationships
models are statistical tools that relate molecular descriptors to a property of interest, such as physicochemical, toxicological, or biological activity. Typically, these models are constructed using machine learning (ML) algorithms. The general workflow for developing models has been extensively documented [2] [3].
Similar to SAR (StructureβActivity Relationship) analysis, modeling requires a dataset of molecules with well-defined structures and experimentally determined activities. Molecular descriptorsβnumerical representations of molecular characteristicsβare calculated for each compound. These descriptors can include atom counts (e.g., number of heteroatoms), fragment counts (e.g., number of benzene rings), topological indices (describing the moleculeβs connectivity), and 3D descriptors (capturing spatial features of the molecules). [4]
@@ II-C Machine Learning Enhancement and Quantum Frontiers
Quantitative Structure-Activity Relationship () models aim to establish a quantitative link between the chemical structure of a molecule and its biological activity [8]. These mathematical models provide an in silico methodology to predict the properties of novel compounds, thereby prioritizing experimental efforts and reducing costs. Applications range from predicting pharmacokinetic processesβabsorption, distribution, metabolism, and excretion ()βto assessing toxicological endpoints, and even modelling new properties.
@@ III Methodology
models are evaluated using various metrics that assess predictive performance [4]. Key metrics include:
β’
Accuracy (ACC)
β’
Precision (PPV)
β’
Sensitivity (Recall, TPR)
β’
Specificity (TNR)
β’
F1-score
β’
Matthews Correlation Coefficient (MCC)
In our research, we use the ROC-AUC as a key metric to benchmark classical and quantum pipelines, as mentioned in the corresponding section.
III-A Dataset Description and Preprocessing
This study is based on a high-quality dataset provided and curated by the team at ProtoQSAR SL, their contribution was fundamental in the selection of compounds, the calculation, and validation of molecular descriptors.
This dataset serves as the foundational benchmark for developing our quantum machine learning model. By establishing a consistent baseline, it enables robust comparisons and benchmarking against other models trained on the same data.
The dataset was compiled from public dataset of (Dual-specificity tyrosine-phosphorylation-regulated kinase 1A) inhibitors [14]. A comprehensive dataset of 1291 molecules exhibiting inhibitory activity against human kinase was systematically compiled from the database (). represents a critical therapeutic target for neurodegenerative disorders, particularly Alzheimerβs disease, due to its role in tau phosphorylation and neuronal dysfunction [15].
@@ III-A1 Molecular Descriptors
NPR1 (Normalized Principal Ratio 1): The ratio PMI1/PMI3, normalized to eliminate molecular size effects. NPR1 values range from 0.5 (spherical) to 1.0 (linear), providing size-independent shape characterization essential for scaffold-hopping applications.
@@
MolLogP (Molecular LogP): Octanol-water partition coefficient calculated using the Wildman-Crippen method implemented in . Values typically range from -3 to +8 for drug-like compounds, with optimal range 1-3 for oral bioavailability.
TPSA (Topological Polar Surface Area): Sum of surfaces of polar atoms (oxygen and nitrogen) in a molecule, calculated using group contributions. TPSA values Β‘140 Γ Β² correlate with good blood-brain barrier permeability.
NumRotatableBonds: Count of rotatable bonds in the molecule, excluding amide bonds. Drug-like molecules typically have 10 rotatable bonds according to Lipinskiβs rule.
@@
BertzCT (Bertz Complexity Index): A graph-theoretical descriptor that quantifies molecular complexity by considering both the number of atoms and their connectivity patterns. Typical values range from 0 for simple molecules (e.g., methane) to over 1000 for complex natural products [16]. Its considers both the number of atoms and their connectivity patterns within a molecular structure. This index helps in understanding and predicting the complexity of molecules, particularly in the context of synthetic planning and analysis of synthetic strategies. The Bertz Complexity Index utilizes graph theory to represent molecules. In this representation, atoms are nodes (vertices) and bonds are edges within the graph.
Chi0n, Chi1n (Kier-Hall Connectivity Indices): Topological indices describing molecular branching and connectivity. Chi0n reflects molecular size, while Chi1n captures first-order connectivity patterns.
@@
NumHBD (Hydrogen Bond Donors): Count of hydrogen bond donor groups (OH, NH). Drug-like molecules typically have 5 donors.
NumHBA (Hydrogen Bond Acceptors): Count of hydrogen bond acceptor atoms (N, O). Optimal range 10 for oral drugs.
FractionCsp3: Fraction of carbons that are sp3 hybridized. Higher Fsp3 values () correlate with increased three-dimensionality and improved developability.
@@ Acknowledgment
This work was supported by the project funded by the Ministerio de Ciencia, InnovaciΓ³n y Universidades, through . We gratefully acknowledge ProtoQSAR SL and the team led by Eva Serrano-Candelas, Laureano E. Carpio, and Rafael Gozalbes for their fundamental contribution in providing, curating, and validating the dataset used in this study.
For more information about this research, please refer to: